OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5 help guys?

OpenStudy (anonymous):

Welcome to OpenStudy!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

First things first, you want to treat sinx like a another variabe (say y), it'll make the problem a lot easier

OpenStudy (anonymous):

It will then be \[\Large 7y^2-14y+2=-5\]

OpenStudy (anonymous):

If you add 5 to each sde, then you'll be left with \[\Large 7y^2-14y+7=0\]

OpenStudy (anonymous):

Alright, can you factor this out now?

OpenStudy (anonymous):

alright, so y^2-2y+1=0

OpenStudy (anonymous):

ok, so now can you factor that out?